# Post-Completion Learning for Language Models

### v1

@@ Introduction
Post-Completion Space Design
@@
Zero-Cost Inference Implementation
@@
Teacher Model for Data Preparation
@@
White-Box Reinforcement Learning Design
@@
Multi-Dimensional Reward Function Design
@@
[…]
@@
[…]
@@
where $R_{\text{pred}}$ is the self-evaluation score output by the model in the reward section, and $R_{\text{true}}$ is the score calculated by the true reward function.
Self-Evaluation Capability Development
@@
Dual-Track SFT Training Strategy
@@
Theoretical Analysis of Post-Completion Learning
@@
Information-Theoretic Foundation and Optimality Analysis
@@
Proof:
Step 1 (Sequential Training Analysis): In sequential training, let $\theta_{SFT}$ be the parameters after SFT and $\theta_{final}$ be the parameters after subsequent RL. The total error can be decomposed as:
@@
[…]
@@
Table 2: Results on GSM8K Dataset (Accuracy)
@@
Tables 2, 3, and 4 present our main experimental results across three reasoning datasets and multiple model scales. PCL demonstrates consistent and substantial improvements over all baseline methods.
@@
Table 3: Results on StrategyQA Dataset (Accuracy)
@@
Table 4: Results on MathQA Dataset (Accuracy)

### v3

@@ Introduction
Post-Completion Space
@@
Inference: Early Stopping
@@
Evaluation Data Preparation
@@
White-Box Reinforcement Learning
@@
Reward Functions
@@
$R_{a}=\mathbf{1}[\text{answer}=\text{ground\_truth}]$ (1)
@@
$R_{f}=\mathbf{1}[\text{contains\_all\_sections}(\text{output})]$
@@
where $R_{\text{pred}}$ is the self-evaluation score output by the model in the reward section, and $R_{\text{true}}$ is the score calculated by the true reward function. Notably, the choice of reward function is not restricted and supports many others based on detailed task.
Self-Evaluation Capability
@@
Dual-Track SFT Strategy
@@
Theoretical Analysis
@@
Information-Theoretic Foundation and Optimality
@@
Proof: Step 1 (Sequential Training Analysis): In sequential training, let $\theta_{SFT}$ be the parameters after SFT and $\theta_{final}$ be the parameters after subsequent RL. The total error can be decomposed as:
@@
Proof: The key insight is that the $\langle\text{post-completion}\rangle$ token creates a natural parameter separation. Let $\Theta_{R}$ and $\Theta_{E}$ denote the parameter sets primarily affected by reasoning and evaluation losses respectively. Due to the sequential attention mechanism, we have $|\Theta_{R}\cap\Theta_{E}|\ll|\Theta_{R}\cup\Theta_{E}|$.
@@
Table 2: Results on GSM8K Dataset (% Accuracy)
@@
Tables 2, 3, and 4 present our main experimental results across three reasoning datasets and multiple model scales. PCL demonstrates consistent and improvements over all baseline methods.
Traditional methods typically employ math_verify to extract answers through regular expression matching on the model’s complete response. In contrast, our approach requires the model to strictly adhere to a predefined format and performs exact matching on the <answer> part. Consequently, our evaluation criteria are relatively more stringent, which may result in lower baseline performance metrics compared to other studies.
@@
Table 3: Results on StrategyQA Dataset (% Accuracy)
@@
Table 4: Results on MathQA Dataset (% Accuracy)